# Beyond Winning Strategies: Admissible and Admissible Winning Strategies for Quantitative Reachability Games

### v2

@@ 1 Introduction
Reactive Synthesis is the problem of automatically generating reactive systems from logical specifications, first proposed by Church [1963]. Its applications span a wide range of domains, including robotics [McMahon et al., 2023; He et al., 2017; Kress-Gazit et al., 2018], program synthesis [Pnueli and Rosner, 1989], distributed systems [Filippidis and Murray, 2016], formal verification [Kupferman and Vardi, 2001], and security [Zhou and Foley, 2003]. Existing approaches to reactive synthesis usually boil down to computing a strategy over a game between a System (Sys) and an Environment (Env) player, with the goal of finding a winning strategy, which guarantees the Sys player achieves its objectives regardless of the Env player’s moves. In quantitative settings, the game incorporates a payoff requirement as part of the Sys player’s objectives. In many scenarios, however, a winning strategy may not exist, resulting in a failure in the synthesis algorithm. Prior works relax this requirement using the notions of best effort [Aminof et al., 2021] and admissibility [Berwanger, 2007; Brandenburger et al., 2008]. Best effort is specific to qualitative games, while admissibility is recently studied in quantitative games under strong assumptions on the Env player, namely, rationality and known objectives [Brenguier et al., 2016]. This paper aims to study admissible strategies without these assumptions in quantitative reachability games with a particular focus on applications in robotics.
@@
To this end, this paper studies admissible strategies [Faella, 2009; Berwanger, 2007] in quantitative reachability games without assumptions on rationality and the objectives of the Env player. These games can model quantitative reactive synthesis for finite-behaviors expressed in, e.g., syntactically co-safe Linear Temporal Logic (ltl) [Kupferman and Vardi, 2001] and ltl over finite behaviors ($\textsc{ltl}_{f}$) [De Giacomo and Vardi, 2013]. We show that admissible strategies relax requirement of winning strategies, and always exist. We prove that, unlike the qualitative setting, quantitative admissible strategies are generally history-dependent even for finite payoff functions. Then, we show that such strategies can produce overly optimistic behaviors, which may be undesirable for robotics applications. To mitigate this, we propose admissible winning strategies, which have the desirable property of enforcing specification satisfaction when possible while being admissible. We prove that similar to admissible strategies, admissible winning strategies always exist and may require finite memory. Then, we provide necessary and sufficient conditions for both strategies and propose synthesis algorithms. Finally, we provide various robotic examples to show that admissible and admissible winning strategies provide desirable and flexible behaviors without a-priori knowledge of the objectives of other agents.
@@
Several works explore alternatives to winning strategies. Specifically, Faella [2009] investigates various concepts in qualitative games with reachability objectives within a zero-sum framework. They focus solely on the Sys player’s objective without making any assumptions about the Env player. They use admissibility to define the notion of best-effort (BE). Aminof et al. [2021]; De Giacomo et al. [2023] further examine the complexity of synthesizing BE strategies, showing that it can be reduced to standard algorithms, with memoryless strategies being sufficient. In contrast, our work considers quantitative reachability games where the objective is to reach a goal state with minimal total cost. We show that memoryless strategies are insufficient in our context, and our synthesis approach does not reduce to standard algorithms.
The notion of admissiblity has also been explored in normal form games [Brandenburger et al., 2008; Apt, 2011]. In qualitative games with logical specifications (extensive form), admissiblity has been investigated for n-player infinite games, where each player has their own objective and is assumed to play admissibly with respective to that objective – referred to as assume admissible (AA). Berwanger [2007] was the first to formalize this notion of AA for games played on graphs. Subsequently, Brenguier et al. [2014, 2015] establish the complexity and give algorithms for $\omega$-regular objectives. In our settings, we consider reachability games that terminate in finite time. Notably, we make no assumptions about the Env player, i.e., we neither know Env player’s objective nor require them to play admissibly.
The work closest to ours is by Brenguier et al. [2016], who study admissibility in quantitative settings. They extend prior work [Brenguier et al., 2014, 2015] on infinite duration qualitative games to quantitative objectives. They give necessary and sufficient conditions for admissible strategies. Unlike their work, in our setting, we consider finite duration games and appropriately define our payoff over finite traces. We show our game is always determined and thus optimal worst-case and cooperative strategies always exist. While our analysis shares some conceptual similarities with theirs, addressing the finite play setting requires a distinct theoretical approach compared to infinite plays. Brenguier et al. [2016] give a sketch of their algorithm based on parity games. We, however, present a detailed yet simpler synthesis algorithm. We also analyze emergent behavior under admissible strategies in robotics settings. We observe that these strategies can be overly optimistic. To address this, we identify the underlying cause and propose the concept of admissible winning strategies to mitigate such optimism.
@@ 2 Problem Formulation
The overarching goal of this work is quantitative reactive synthesis for $\textsc{ltl}_{f}$ or cosafe ltl specifications where satisfaction cannot necessarily be guaranteed. This problem reduces to reachability analysis in quantitative games played between the Sys and Env players [Baier and Katoen, 2008]. For the sake of generality, we focus on these games.
2-Player Quantitative Games and Strategies
@@
The players choose actions according to a strategy.
@@
Given strategies $\sigma$ and $\tau$, a unique sequence of states, called play and denoted by $P^{v_{0}}(\sigma,\tau)$, is induced from $v_{0}$. A play can be finite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots v_{n}\in V^{*}$ or infinite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots\in V^{\omega}$.
We denote by $\operatorname{Plays}^{v}:=\{P^{v}(\sigma,\tau)\mid\sigma\in\Sigma,\;\tau\in\mathrm{T}\}$ the set of plays starting from $v$ under every Sys and Env strategy. $\operatorname{Plays}^{v}(\sigma)$ is the set of plays induced by a fixed strategy $\sigma$ and every Env strategy. We note that a finite play occurs iff a goal state is reached.
@@
Given $h$, Sys ($\sigma$) and Env ($\tau$) strategies compatible with $h$, let $P^{h}(\sigma,\tau)$ denote a play that extends history $h$ The cooperative value $\operatorname{cVal}(h,\sigma)$ is the payoff of the play such that the Env player plays minimally, i.e.,
@@
Beyond Winning Strategies
In many applications, when a winning strategy does not exist, it is desirable for the Sys player to adopt a strategy that ensures the best possible outcome. For game $\mathcal{G}$ in Fig. 1(b), $\sigma_{win}$ does not exist but it is better for the Sys player to keep trying, i.e., move to $v_{1}$ or $v_{3}$ rather than giving up. To capture this intuition, we turn to the classical notion of dominance.
@@
Definition 5 (Dominance [Brenguier et al., 2016; Berwanger, 2007]).
@@
In the game in Fig. 2, let $\sigma_{1}:(v_{0}\to v_{1}),(v_{4}\to v_{5})$, $\sigma_{2}:(v_{0}\to v_{1}),(v_{4}\to v_{7})$, and $\sigma_{3}:(v_{0}\to v_{2}),(v_{3}\to v_{2})$. As $\sigma_{2}\succ\sigma_{1}$, $\sigma_{1}$ is not admissible. But, $\sigma_{3}$ is not weakly dominated by $\sigma_{2}$ as there exists a play under which $\sigma_{3}$ does strictly better than $\sigma_{2}$ for every play in $\operatorname{Plays}^{v_{0}}(\sigma_{2})$. Hence, $\sigma_{3}$ and $\sigma_{2}$ are both admissible.
Brandenburger et al. [2008] rationalize a player playing admissibly to be doing their best. Thus, the first problem we consider is the synthesis of admissible strategies.
@@
Intuitively, $\sigma\in\Sigma_{adm}$ allows the Sys player to do its best without any assumptions about the Env player. In qualitative settings, winning strategies are always admissible [Brenguier et al., 2015, Lemma 8]. In quantitative settings, however, winning strategies are not necessarily admissible (e.g., $\sigma_{1}$ in Fig. 2). Thus, to compute enforceable admissible strategies, we introduce the notion of an admissible winning strategy.
@@ 3 Admissible Strategies
Admissible Strategies are not Value-Preserving
@@
Unlike the qualitative setting [Faella, 2009; Aminof et al., 2020], Lemma 1 shows that we cannot characterize admissible strategies solely on the basis of $\operatorname{\mathbf{WCO}}$ strategies as they are not value-preserving in our quantitative setting. Below, we derive two new categories of strategies from $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ that are always admissible. We then discuss their properties and show that they are also necessary conditions for admissibility. Table 1 summarizes properties of all new strategies we define hereafter.
@@
Characterization of Admissible Strategies
@@
From Lemma 3, it suffices for us to show that $\operatorname{\mathbf{WCo-Op}}$ strategies always exist for admissible strategies to always exist. Interestingly, unlike prior work [Brenguier et al., 2016], in our case, $\operatorname{\mathbf{WCo-Op}}$ strategies always exist and at least one is memoryless. The latter is desirable because memoryless strategies can be computed efficiently using fixed-point-based algorithms [Baier and Katoen, 2008].
@@
Existence of Admissible Strategies
@@
The proof follows from Lemma 3 and Theorem 2. Observe that Theorem 3 characterizes the set of all admissible strategies and Theorem 4 establishes their existence in full generality, i.e., independent of $\mathcal{B}$, for 2-player turn-based games with reachability objectives. For computability considerations, we bound the payoffs associated with plays to a given budget $\mathcal{B}$ so that the set of all admissible strategies is finite. This is a reasonable assumption that is often used in, e.g., energy games with fixed initial credit and robotics application with finite resources [Chakrabarti et al., 2003; Bouyer et al., 2008; He et al., 2017; Filiot et al., 2010].
Admissible Strategy Synthesis
@@ 11                   $\Sigma_{adm}:h\to v^{\prime}$ Add only states in $h$
Game Tree Arena
@@
After constructing $\mathcal{G}^{\prime}$, Alg. 1 computes $\operatorname{aVal}$ and $\operatorname{cVal}$ values for each state in $\mathcal{G}^{\prime}$ using the Value Iteration algorithm [Brihaye et al., 2017]. If $\mathcal{B}<\operatorname{cVal}(v_{0})$, then there does not exist a play that reaches a goal state in $\mathcal{G}^{\prime}$. Thus, all strategies in $\mathcal{G}$ are admissible. To compute admissible strategies on $\mathcal{G}^{\prime}$, we use a DFS algorithm to traverse every play and check if the admissibility criteria from Lemma 4 is satisfied. If yes, we add the history and the successor state to $\Sigma_{adm}$. We repeat this until every state in $\mathcal{G}^{\prime}$ is explored. This algorithm is sound and complete with polynomial time complexity.
@@
Given $\mathcal{G}$ and a fixed budget $\mathcal{B}$, Alg. 1 returns the set of all admissible strategies $\Sigma_{adm}$ and has polynomial time complexity.
@@ 4 Admissible Winning Strategies
Optimistic Strategy
In Sec. 3, we show that $\operatorname{\mathbf{SC}}$ strategies are willing to risk a higher payoff $(\operatorname{aVal}(\delta(\operatorname{last}(h),\sigma(h)))>\operatorname{cVal}(h,\sigma))$ in the hopes that the Env will cooperate, i.e., they are optimistic.
@@
Admissible Winning Strategy Synthesis
@@
The algorithm is similar to Alg. 1 except for the admissibility checking criteria. The tree construction, syntax, and semantics are as outlined in Sec. 3. We modify admissibility checking criteria in Line 9 of Alg. 1 to $\operatorname{cVal}(v^{\prime})<\min\{\operatorname{aValues}\}$ and $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ This modification expresses the value-preserving term from Eq. (9a).
@@
Given $\mathcal{G}$ and a fixed budget $\mathcal{B}$, the algorithm described returns the set of all admissible winning strategies $\Sigma^{win}_{adm}$.
@@ 5 Illustrative Examples
We now discuss the emergent behavior under admissible and admissible winning strategies in two settings: (i) a gridworld domain, where two agents take actions in turns, and (ii) a manipulator domain, where a robotic arm and a human operate in a shared workspace. For both domains, the task $\varphi$ is specified using $\textsc{ltl}_{f}$ formulas. We first construct a game abstraction $\mathcal{G}$. Next, we construct a deterministic finite automaton (DFA) for the task [Fuggitti, 2019] and take the product to construct the product (DFA) game, where the objective for the Sys player is to reach a set of goal (accepting) states. Code is included in the supplementary material and will be made public upon publication. Additionally, more manipulator domain experiments and empirical validations of PTIME complexity of Alg. 1 for fixed $\mathcal{B}$ are provided in the Appendix. Videos for all experiments are provided in the supplementary material.
@@
Fig. 3 illustrates a gridworld domain with Sys (red), Env (blue), goal (green), and lava (orange) states. The objective for Sys is to reach the goal state while avoiding the Env player. The players should not enter lava and Env player cannot traverse through goal state. Action cost is 1 for all Sys player actions. The game is played in turns starting with Sys where both players take a step in each cardinal direction. Figs. 3(a) and 3(c) illustrate the initial position of both players for two different scenarios. Figs. 3(a)-3(b) and 3(c)-3(d) illustrate $\sigma_{adm}$ and $\sigma^{win}_{adm}$ for different budgets. In Figs. 3(a)-3(b), a winning strategy exists if $\mathcal{B}\geq 12$. Hence, for $\mathcal{B}=5$ (Fig. 3(a)), only $\sigma_{adm}$ exists, which relies on Env’s cooperation to reach goal state. For $B=12$ (Fig. 3(b)), $\sigma^{win}_{adm}$ exists and commits to go around and reach the goal state. Note that $\sigma_{adm}$ (not shown) also exists but does not commit to go around.
@@
Comparison against Best-Effort (BE): Here we illustrate differences between BE and admissible strategies. Besides qualitative vs quantitative, in the winning region, a major difference is that every $\sigma_{win}$ is also BE, but not every $\sigma_{win}$ is winning admissible. In Fig. 3(d), if Env player stays in the left corner (cooperative), $\sigma^{win}_{adm}$ commits to dashed-line strategy as the solid-line strategy is not $\operatorname{\mathbf{Co-Op}}$, whereas BE does not distinguish between the two strategies as both are winning.
Figure 4: Left: Initial setup. Right: $\sigma_{adm}$ for $\mathcal{B}=13$. Task: $\varphi=\bigl{(}p_{04}\wedge p_{15}\wedge\raisebox{-1.04625pt}{\LARGE$\circ$}(p_{26}\vee p_{36})\bigr{)}$ where $p_{ij}$ is block $i$ at loc $j$. Blocks 0 to 2 are boxes and block 3 is a cylinder.
@@
This domain considers a robotic arm (Sys) operating in presence of a human (Env), as described in [Muvvala et al., 2022]. The task for the robot is to build an arch with two boxes as support and one block (cylinder or box) on top as shown in Fig. 4. The human can choose to either move a block back to the initial position or not intervene. The robot’s actions are transit, grasp, transfer, and release with unit cost for each action, except for transferring the cylinder, which costs two units.

### v3

@@ 1 Introduction
Reactive Synthesis is the problem of automatically generating reactive systems from logical specifications, first proposed by Church (1963). Its applications span a wide range of domains, including robotics McMahon et al. (2023); He et al. (2017); Kress-Gazit et al. (2018), program synthesis Pnueli and Rosner (1989), distributed systems Filippidis and Murray (2016), formal verification Kupferman and Vardi (2001), and security Zhou and Foley (2003). Existing approaches to reactive synthesis usually boil down to computing a strategy over a game between a System (Sys) and an Environment (Env) player, with the goal of finding a winning strategy, which guarantees the Sys player achieves its objectives regardless of the Env player’s moves. In quantitative settings, the game incorporates a payoff requirement as part of the Sys player’s objectives. In many scenarios, however, a winning strategy may not exist, resulting in a failure in the synthesis algorithm. Prior works relax this requirement using the notions of best effort Aminof et al. (2021) and admissibility Berwanger (2007); Brandenburger et al. (2008). Best effort is specific to qualitative games, while admissibility is recently studied in quantitative games under strong assumptions on the Env player, namely, rationality and known objectives Brenguier et al. (2016). This paper aims to study admissible strategies without these assumptions in quantitative reachability games with a particular focus on applications in robotics.
@@
To this end, this paper studies admissible strategies Faella (2009); Berwanger (2007) in quantitative reachability games without assumptions on rationality and the objectives of the Env player. These games can model quantitative reactive synthesis for finite-behaviors expressed in, e.g., syntactically co-safe Linear Temporal Logic (ltl) Kupferman and Vardi (2001) and ltl over finite behaviors ($\textsc{ltl}_{f}$) De Giacomo and Vardi (2013). We show that admissible strategies relax requirement of winning strategies, and always exist. We prove that, unlike the qualitative setting, quantitative admissible strategies are generally history-dependent even for finite payoff functions. Then, we show that such strategies can produce overly optimistic behaviors, which may be undesirable for robotics applications. To mitigate this, we propose admissible winning strategies, which have the desirable property of enforcing specification satisfaction when possible while being admissible. We prove that similar to admissible strategies, admissible winning strategies always exist and may require finite memory. Then, we provide necessary and sufficient conditions for both strategies and propose synthesis algorithms. Finally, we provide various robotic examples to show that admissible and admissible winning strategies provide desirable and flexible behaviors without a-priori knowledge of the objectives of other agents.
@@
Several works explore alternatives to winning strategies. Specifically, Faella investigates various concepts in qualitative games with reachability objectives within a zero-sum framework. They focus solely on the Sys player’s objective without making any assumptions about the Env player. They use admissibility to define the notion of best-effort (BE). Aminof et al.; De Giacomo et al. further examine the complexity of synthesizing BE strategies, showing that it can be reduced to standard algorithms, with memoryless strategies being sufficient. In contrast, our work considers quantitative reachability games where the objective is to reach a goal state with minimal total cost. We show that memoryless strategies are insufficient in our context, and our synthesis approach does not reduce to standard algorithms.
The notion of admissiblity has also been explored in normal form games Brandenburger et al. (2008); Apt (2011). In qualitative games with logical specifications (extensive form), admissiblity has been investigated for n-player infinite games, where each player has their own objective and is assumed to play admissibly with respective to that objective – referred to as assume admissible (AA). Berwanger was the first to formalize this notion of AA for games played on graphs. Subsequently, Brenguier et al.; Brenguier et al. establish the complexity and give algorithms for $\omega$-regular objectives. In our settings, we consider reachability games that terminate in finite time. Notably, we make no assumptions about the Env player, i.e., we neither know Env player’s objective nor require them to play admissibly.
The work closest to ours is by Brenguier et al. (2016), who study admissibility in quantitative settings. They extend prior work Brenguier et al. (2014, 2015) on infinite duration qualitative games to quantitative objectives. They give necessary and sufficient conditions for admissible strategies. Unlike their work, in our setting, we consider finite duration games and appropriately define our payoff over finite traces. We show our game is always determined and thus optimal worst-case and cooperative strategies always exist. While our analysis shares some conceptual similarities with theirs, addressing the finite play setting requires a distinct theoretical approach compared to infinite plays. Brenguier et al. give a sketch of their algorithm based on parity games. We, however, present a detailed yet simpler synthesis algorithm. We also analyze emergent behavior under admissible strategies in robotics settings. We observe that these strategies can be overly optimistic. To address this, we identify the underlying cause and propose the concept of admissible winning strategies to mitigate such optimism.
@@ 2 Problem Formulation
The overarching goal of this work is quantitative reactive synthesis for $\textsc{ltl}_{f}$ or cosafe ltl specifications where satisfaction cannot necessarily be guaranteed. This problem reduces to reachability analysis in quantitative games played between the Sys and Env players Baier and Katoen (2008). For the sake of generality, we focus on these games.
2.1 2-Player Quantitative Games and Strategies
@@
The players choose actions according to a strategy. Formally,
@@
Given strategies $\sigma$ and $\tau$, a unique sequence of states, called play and denoted by $P^{v_{0}}(\sigma,\tau)$, is induced from $v_{0}$. Note that the play is unique for terminating plays only. A play can be finite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots v_{n}\in V^{*}$ or infinite $P^{v_{0}}(\sigma,\tau):=v_{0}v_{1}\dots\in V^{\omega}$. We denote by $\operatorname{Plays}^{v}:=\{P^{v}(\sigma,\tau)\mid\sigma\in\Sigma,\;\tau\in\mathrm{T}\}$ the set of plays starting from $v$ under every Sys and Env strategy. $\operatorname{Plays}^{v}(\sigma)$ is the set of plays induced by a fixed strategy $\sigma$ and every Env strategy. We note that a finite play occurs iff a goal state is reached.
@@
Given $h$, Sys ($\sigma$) and Env ($\tau$) strategies compatible with $h$, let $P^{h}(\sigma,\tau)$ denote a play that extends history $h$. The cooperative value $\operatorname{cVal}(h,\sigma)$ is the payoff of the play such that the Env player plays minimally, i.e.,
@@
2.2 Beyond Winning Strategies
In many applications, when a winning strategy does not exist, it is desirable for the Sys player to adopt a strategy that ensures the best possible outcome. For game $\mathcal{G}$ in Fig. 1(b), $\sigma_{win}$ does not exist but it is better for the Sys player to keep trying, i.e., move to $v_{1}$ or $v_{3}$ rather than giving up. To capture this intuition, we turn to the classical notion of dominance Leyton-Brown and Shoham (2008).
@@ 2.2 Beyond Winning Strategies
Definition 5 (Dominance Brenguier et al. (2016); Berwanger (2007)).
@@
In the game in Fig. 2, let $\sigma_{1}:(v_{0}\to v_{1}),(v_{4}\to v_{5})$, $\sigma_{2}:(v_{0}\to v_{1}),(v_{4}\to v_{7})$, and $\sigma_{3}:(v_{0}\to v_{2}),(v_{3}\to v_{2})$. As $\sigma_{2}\succ\sigma_{1}$, $\sigma_{1}$ is not admissible. But, $\sigma_{3}$ is not weakly dominated by $\sigma_{2}$ as there exists a play under which $\sigma_{3}$ does strictly better than $\sigma_{2}$ for every play in $\operatorname{Plays}^{v_{0}}(\sigma_{2})$. Hence, $\sigma_{3}$ and $\sigma_{2}$ are both admissible. This example demonstrates that (i) admissible strategies can be overly optimistic, and (ii) they differ from winning strategies.
Brandenburger et al. rationalize a player playing admissibly to be doing their best. Thus, the first problem we consider is the synthesis of admissible strategies.
@@
Intuitively, $\sigma\in\Sigma_{adm}$ allows the Sys player to do its best without any assumptions about the Env player. In qualitative settings, winning strategies are always admissible (Brenguier et al., 2015, Lemma 8). In quantitative settings, however, winning strategies are not necessarily admissible (e.g., $\sigma_{1}$ in Fig. 2). Thus, to compute enforceable admissible strategies, we introduce the notion of an admissible winning strategy.
@@ 3 Admissible Strategies
3.1 Admissible Strategies are not Value-Preserving
@@
Unlike the qualitative setting Faella (2009); Aminof et al. (2020), Lemma 1 shows that we cannot characterize admissible strategies solely on the basis of $\operatorname{\mathbf{WCO}}$ strategies as they are not value-preserving in our quantitative setting. Below, we derive two new categories of strategies from $\operatorname{\mathbf{WCO}}$ and $\operatorname{\mathbf{Co-Op}}$ that are always admissible. We then discuss their properties and show that they are also necessary conditions for admissibility. Table 1 summarizes properties of all new strategies we define hereafter.
@@
3.2 Characterization of Admissible Strategies
@@
From Lemma 3, it suffices for us to show that $\operatorname{\mathbf{WCo-Op}}$ strategies always exist for admissible strategies to always exist. Interestingly, unlike prior work Brenguier et al. (2016), in our case, $\operatorname{\mathbf{WCo-Op}}$ strategies always exist and at least one is memoryless. The latter is desirable because memoryless strategies can be computed efficiently using fixed-point-based algorithms Baier and Katoen (2008).
@@
3.3 Existence of Admissible Strategies
@@
The proof follows from Lemma 3 and Theorem 2. Observe that Theorem 3 characterizes the set of all admissible strategies and Theorem 4 establishes their existence in full generality, i.e., independent of $\mathcal{B}$, for 2-player turn-based games with reachability objectives. For computability considerations, we bound the payoffs associated with plays to a given budget $\mathcal{B}$ so that the set of all admissible strategies is finite. This is a reasonable assumption that is often used in, e.g., energy games with fixed initial credit and robotics application with finite resources Chakrabarti et al. (2003); Bouyer et al. (2008); Muvvala and Lahijanian (2023); Filiot et al. (2010).
3.4 Admissible Strategy Synthesis
@@ 11                   $\Sigma_{adm}:h\to v^{\prime}$ Add only states in $h$
Game Tree Arena.
@@
After constructing $\mathcal{G}^{\prime}$, Alg. 1 computes $\operatorname{aVal}$ and $\operatorname{cVal}$ values for each state in $\mathcal{G}^{\prime}$ using the Value Iteration algorithm Brihaye et al. (2017). If $\mathcal{B}<\operatorname{cVal}(v_{0})$, then there does not exist a play that reaches a goal state in $\mathcal{G}^{\prime}$. Thus, all strategies in $\mathcal{G}$ are admissible. To compute admissible strategies on $\mathcal{G}^{\prime}$, we use a DFS algorithm to traverse every play and check if the admissibility criteria from Lemma 4 is satisfied. If yes, we add the history and the successor state to $\Sigma_{adm}$. We repeat this until every state in $\mathcal{G}^{\prime}$ is explored. This algorithm is sound and complete with polynomial time complexity.
@@
Given $\mathcal{G}$ and a budget $\mathcal{B}$, Alg. 1 returns the set of all admissible strategies $\Sigma_{adm}$. The algorithm runs in polynomial time when $\mathcal{B}$ is fixed and in pseudo-polynomial time when $\mathcal{B}$ is arbitrary.
@@ 4 Admissible Winning Strategies
4.1 Optimistic Strategy
In Sec. 3.1, we show that $\operatorname{\mathbf{SC}}$ strategies are willing to risk a higher payoff $(\operatorname{aVal}(\delta(\operatorname{last}(h),\sigma(h)))>\operatorname{cVal}(h,\sigma))$ in the hopes that the Env will cooperate, i.e., they are optimistic.
@@ 4.1 Optimistic Strategy
4.2 Admissible Winning Strategy Synthesis
@@
The algorithm is similar to Alg. 1 except for the admissibility checking criteria. The tree construction, syntax, and semantics are as outlined in Sec. 3.4. We modify admissibility checking criteria in Line 9 of Alg. 1 to $\operatorname{cVal}(v^{\prime})<\min\{\operatorname{aValues}\}$ and $\neg(v\in V_{win})\vee(v^{\prime}\in V_{win})$ This modification expresses the value-preserving term from Eq. (9a).
@@
Given $\mathcal{G}$ and budget $\mathcal{B}$, the algorithm described returns the set of all admissible winning strategies $\Sigma^{win}_{adm}$ and has same time complexity as Alg. 1.
@@ 5 Illustrative Examples
We now discuss the emergent behavior under admissible and admissible winning strategies in two settings: (i) a gridworld domain, where two agents take actions in turns, and (ii) a manipulator domain, where a robotic arm and a human operate in a shared workspace. For both domains, the task $\varphi$ is specified using $\textsc{ltl}_{f}$ formulas. We first construct a game abstraction $\mathcal{G}$. Next, we construct a Deterministic Finite Automaton for the task Fuggitti (2019) and take the product to construct the product game, where the objective for the Sys player is to reach a set of goal (accepting) states. Additional experiments and empirical validations of PTIME complexity of Alg. 1 for fixed $\mathcal{B}$ are provided in the extended version Muvvala et al. (2025). Code is available on Github Muvvala (2025).
@@
Fig. 3 illustrates a gridworld domain with Sys (red), Env (blue), goal (green), and lava (orange) states. The objective for Sys is to reach the goal state while avoiding the Env player. The players should not enter lava and Env player cannot traverse through goal state. Action cost is 1 for all Sys player actions. The game is played in turns starting with Sys where both players take a step in each cardinal direction.
Figs. 3(a) and 3(c) illustrate the initial position of both players for two different scenarios. Figs. 3(a)-3(b) and 3(c)-3(d) illustrate $\sigma_{adm}$ and $\sigma^{win}_{adm}$ for different budgets. In Figs. 3(a)-3(b), a winning strategy exists if $\mathcal{B}\geq 12$. Hence, for $\mathcal{B}=5$ (Fig. 3(a)), only $\sigma_{adm}$ exists, which relies on Env’s cooperation to reach goal state. For $B=12$ (Fig. 3(b)), $\sigma^{win}_{adm}$ exists and commits to go around and reach the goal state. Note that $\sigma_{adm}$ (not shown) also exists but does not commit to go around.
@@
Comparison against Best-Effort (BE): Here we illustrate differences between BE and admissible strategies. Besides qualitative vs quantitative, in the winning region, a major difference is that every $\sigma_{win}$ is also BE, but not every $\sigma_{win}$ is admissible winning. In Fig. 3(d), if Env player stays in the left corner (cooperative), $\sigma^{win}_{adm}$ commits to dashed-line strategy as the solid-line strategy is not $\operatorname{\mathbf{Co-Op}}$, whereas BE does not distinguish between the two strategies as both are winning.
Figure 4: Left: Initial setup. Right: $\sigma_{adm}$ for $\mathcal{B}=13$. Task: $\varphi=\bigl{(}p_{04}\wedge p_{15}\wedge\raisebox{-1.04625pt}{\LARGE$\circ$}(p_{26}\vee p_{36})\bigr{)}$ where $p_{ij}$ is block $i$ at loc $j$. Blocks 0 to 2 are boxes and block 3 is a cylinder. Supplementary Video: https://youtu.be/t0TMOC_PrNk
@@
This domain considers a robotic arm (Sys) operating in presence of a human (Env), as described in Muvvala et al. (2022). The task for the robot is to build an arch with two boxes as support and one block (cylinder or box) on top as shown in Fig. 4. The human can choose to either move a block back to the initial position or not intervene. The robot’s actions are transit, grasp, transfer, and release with unit cost for each action, except for transferring the cylinder, which costs two units.